# DIVERSE: A Dataset of YouTube Video Comment Stances with a Data Programming Model

### v2

@@ Abstract
Stance detection of social media text is a key component of many real-world applications like evaluating marketing campaigns, evaluating political policies or candidates, or evaluating information environments. However, creating automatic stance labeling systems requires the manual annotation of stances, which is both tedious and resource-intensive. This paper introduces an stance labeling method that makes use of weak signals of sentence tone, then consolidating these signals with a Data Programmingmodel for the final stance label. In a time of international conflict, understanding the public opinion towards the country’s military is crucial for recruitment. We present DIVERSE, a dataset involve stances towards YouTube videos of the US military [^1] Datasetavailableathttps://doi.org/10.5281/zenodo.10493803. On average, the videos have 200 comments each, and the stances skew slightly towards the “against” characterization for both the US army and the video.
@@ I Introduction
Today’s world is rife with armed conflict: the Russia-Ukraine war, the Israel-Hamas conflict, the Yemeni Civil War, Somali Civil War… and so forth. The US is involved in a number of these conflicts, and also has the mammoth task of defending her own. During conflict, it is crucial to know the stance of the people towards its own defending forces, as a gauge of the support the forces have from the public. However, US military is facing a recruitment crisis. For the last few years, the US Army have struggled to meet recruiting targets to maintain their desired force structure (Brown, 2023). Social media plays an important role in discerning stances (AlDayel and Magdy, 2021), and in recruitment (Ng and Cruickshank, 2023). Thus, it is critical to understand how messages from the military is received on social media, for it may be used to shape beliefs of users. More broadly speaking, the ability to evaluate how social media content is being received by a target audience is crucial for any marketing campaign, be it a war recruiting effort or a organizational recruitment effort or for influencing consumer behavior.
@@
A particular indicator of the reception of social media content is stance. Stance detection, or the determination of an expressed or implied opinion toward a target, remains a primary task for many computational social studies. Several articles have proposed methods and datasets for the task of stance detection (Allaway and McKeown, 2023; Du Bois, 2007). However, existing datasets explore stance as expressed on a single social media platform, mainly X, and only have a single stance per comment (Sobhani et al., 2017; Conforti et al., 2020; Mohammad et al., 2016; Hossain et al., 2020; Kochkina et al., 2018; Villa-Cox et al., 2020). Since stance detection remains a difficult task due to the context-dependent nature of the definition of stance (e.g., the requirement of a target for stance to be defined), along with varying definitions of stance (Ng and Carley, 2022), all of the aforementioned stance datasets rely on human annotation to create the stance labels. This can result in human errors being present in the datasets as well as greatly reduce the size of the datasets to be in the thousands since human annotation requires a lot of resources and time.
@@
To address these shortcomings in prior stance dataset, we propose a new stance dataset based on YouTube comments and labeled with machine labeling through the data programming paradigm of weak supervision. This dataset is a collection of YouTube comments posted on videos uploaded by the US Army [^2] Offical US Army YouTube Channel: https://www.youtube.com/@usarmy. The comments are annotated for weak labels representing tone (hate speech, sarcasm, sentiment), inference of two LLMs (stance towards video, stance towards army), then combining these labels to produce the final stance labels of the comment towards the US Army. With this dataset, we develop a automated and systematic approach to labeling stances, aiding in understanding critical issues with the public opinion towards the army and its recruitment efforts.
Our contributions of this paper are as follows:
  • The use of an data-programming based stance-labeling methodology. We combine recent advances for creating large, high-quality labeled datasets Ratner et al. (2016, 2019) with contextual knowledge to generate quality stance labels. This involves making use of weak labels such as the presence of hate speech, the sentiment of the content, the presence of sarcasm, known keywords, and the evaluation of Large Language Models (LLMs), which are combined using weak supervision to determine the final stance label.
@@
  • A novel benchmark stance dataset of YouTube comments. We present a dataset for the task of stance detection that has several novel differences from previous benchmark datasets. While most stance datasets use posts from Twitter, ours provides posts from the YouTube platform instead, contributing to the breadth of datasets. Another difference is that our dataset labels stances towards more than one target in a comment (the only other being Sobhani et al. (2017)), and provides labels to weak signals including: the presence of contentious issues, conspiracy theories, and other emotional content. Finally, this dataset has more than three times the number of labeled comments over the current largest benchmark dataset (Conforti et al. (2020) at ${}32,000$ data points vs 173,013 data points).
@@ III Dataset Creation
This dataset contains comments of YouTube comments, each labeled for a set of two final stances: (a) the stance of the comment towards its source video, and (b) the stance of the comment towards the US Army. There are three possible options for stances: “supports”, “against”, and “neutral”. This dataset also contains the weak labels used to fold into the final stance through weak supervision.
@@ III-A Data Collection
This dataset was collected to understand the reception of YouTube users toward the US Army’s marketing material on YouTube, thereby understanding their support/opposition towards the military. We collected comments on the videos of the US Army’s official YouTube channel (username: @usarmy [^3] https://www.youtube.com/usarmy) using the YouTube Data API [^4] https://developers.google.com/youtube/v3/docs/comments/list. This returned both comments as well as replies to comments. Collection took place from 2 to 5 October 2023, where we retrieved a total of 173,013 comments across 858 videos. For each comment, we collected the following information:
@@
  • video_id: the unique id of the video the comment is commenting on
  • name: the title of the video the comment is commenting on
  • description: a text description of the video the comment is commenting on
@@ III-B Data Annotation Methodology
Instead of using human annotators for stance labeling, we used weak supervision Ratner et al. (2016, 2019). This paradigm produces a final, more accurate label using a series of lesser quality labels, weak labels, which may capture different aspects of the concept being labeled. Such labeling techniques have been employed in labeling happiness with cognitive appraisal dimension Liu and Jaidka (2023) and negotiation strategies using the presence of reasoning and friendliness Jaidka et al. (2023). We used a combination of weak labels based upon subject matter expertise, domain expertise, machine learning models, and recent advances in the use of LLMs for stance detection. These labels were then combined through a weak supervision model to produce the final labels. Figure 1, depicts the entire data annotation methodology workflow. We detail key aspects of the workflow in the following sections.
Figure 1: Pipeline of stance labeling methodology note: change LLM box to 2 boxes not 3
@@ III-C Annotating Weak Labels
To label the comments for stance, we first produced weak labels. To do this we leveraged dataset-specific traits, domain expertise and LLMs. For example, in this dataset, the presence of hate speech or sarcasm frequently indicates the comment being against the video and by extension, the user that uploads it, which is the US Army. Therefore, the presence of these weak signals can be combined to indicate the final label. It is important to note for any given weak label, it need not produce a label for every comment in the dataset. Some of the weak labels simply do not apply to some of the comments, in which case they have an “abstain” response, which is distinct from the possible stance labels, and are not used in the computation of the final stance labels.
@@ III-C1 Hate Speech
Hate Speech is defined as speech to incite violence or hate, and typically has specific targets like ethnic groups or religion Fortuna and Nunes (2018). In the case of this dataset and stance, we noticed that many of the comments against both the posted videos and the US Army more generally featured aspects of hate speech, like antisemitic and misogynistic content. Several machine learning algorithms ranging from logistic regression to neural network models have been constructed to perform automatic hate speech detection Fortuna and Nunes (2018). For this weak label, we used the dehatebert model from HuggingFace [^5] Hate-speech-CNERG/dehatebert-mono-english Aluru et al. (2020), a multi-lingual fine-tuned BERT model. This model was used as is without any fine-tuning. For each input text, this model returns a binary label of “hate” or “no_hate”, and we considered the “hate” response as an “against” stance label.
@@ III-C2 Sarcasm
Sarcasm is a literary device to convey one’s intention in an implicit manner Chaudhari and Chandankhede (2017). As with hate speech, we found that commenters on these YouTube videos often use sarcasm as a way of attacking the video or the entity behind the video. We used a multilingual sarcasm detector666helinivan/multilingual-sarcasm-detector which trains a BERT-based sarcasm detector from newspapers text in English, Dutch and Italian. For each input text, this model returns a label of “sarcasm” or “no_sarcasm”, and we considered the “sarcasm” response as an “against” stance label
@@ III-C3 Sentiment
Sentiment analysis provides an understanding of the positivity or negativity of public sentiment towards a topic Ibrohim et al. (2023). In the context of stance detection and classification, while stance is distinct from sentiment, the two often overlap; a negative sentiment often indicates a stance against a target and vice-versa for a positive sentiment. We used a multilingual-cased sentiment analysis model777lxyuan/distilbert-base-multilingual-cased-sentiments-student that performs sentiment-analysis using a teacher-student architecture with the DistillBERT model. For each input text, this model returns a score between 0 and 1 for three possible sentiment labels of “positive”, “negative”, and “neutral”. We considered a “positive” sentiment greater than 0.75 as a “supports” stance and a “negative” sentiment greater than 0.75 as a “against” stance and any other value as a “neutral” stance.
@@ III-C4 Use of Keywords
We manually determined a set of keywords and phrases that upon present, would represent either one of the stance labels. For each comment, we identified if any words from the set of keywords were present and assigned the corresponding stance to this weak label. For example, keywords and phrases like “hooah”, “awesome”, and “god bless” were associated with a “supports” stance, while “woke”, “murder”, and “propaganda” were associated with an “against” stance.
@@ III-C5 Stance from Large Language Models
Due to the size of the dataset, we opted to use open-source LLMs which could be run on local hardware. Specifically, we used two different LLMs: (a) Flan-UL2 (Tay et al., 2022), which is an encoder-decoder architecture and (b) Mistral-7B (Jiang et al., 2023), which is a decoder-only architecture. We used the Flan-UL2 model in a zero-shot setting (i.e., we did not perform any fine-tuning on the model) (Ziems et al., 2023). For the Mistral-7B model, we fine-tuned the model on a collection of benchmark stance datasets, including Semeval2016 (Mohammad et al., 2016), Election2016 (Sobhani et al., 2017), srq (Villa-Cox et al., 2020), wtwt (Conforti et al., 2020), phemerumors (Kochkina et al., 2018), and covid-lies (Hossain et al., 2020). Based on the research in (Ng and Carley, 2022), which found that the inclusion of more benchmark datasets into training improved the generalizability of the trained model, hence we trained on all of the datasets. For the training of the Mistral LLM, we used a LoRA adapter (Hu et al., 2021) with parameters of $\alpha=1$, $r=16$, $dropout=0.1$, and no bias term. The model was trained for two epochs using an AdamW optimizer (Loshchilov and Hutter, 2017), a learning rate of $41e-$ 4, and a linear learning rate scheduler with 50 warm-up steps. To feed the data to the model during training, we adopted a prompt, with two variations in the task portion of the prompt, that included context about stance classification, as has been done in zero-shot prompting for stance in other works. The following was the training prompt:
@@
Where the [dataset entity] was varied depending on the benchmark dataset (e.g., for Election2016, it was ‘politician’, while for phemerumors it was ‘rumor’). For the other variation of the prompt, the “Analyze…” sentence was replaced with a question-based tasking in the form of: “What is the stance of the following social media statement towards the provided [dataset entity]”. For the srq dataset, we changed the prompt slightly to incorporate the post that the reply is replying to as the task in that dataset is to label the stance of the reply to the post, versus the stance of the post about a topic or person. The loss was determined by whether the model could output the tokens for the correct stance, and only those tokens. The training of the model was done on a local computer running Ubuntu Linux with a 96-core processor and three A6000 GPUs. Overall, this resulted in three, different LLMs for labeling the comments: one off-the-shelf UL2 model and two, fine-tuned Mistral models.
@@
In analyzing the comments for stances using LLMs, we used two different prompts, one that only included the comment, and one that included the comment that a comment was replying to, if it was replying to a comment, with two different variations on each prompt. The labeling prompt without incorporating the reply structure is:
@@
As with the training prompt, the variations for each prompt came from changing the task part of the prompt (i.e., the “Analyze…”) sentence to a question-based task prompt. These variations were only used with the Mistral model fine-tuned by the question prompt Thus, each comment was labeled for a stance by an LLM six different times: once with the reply prompt and once without the reply prompt, for the UL2 model, the stance-tuned Mistral model, and the stance-tuned Mistral model with the question-task prompt. Since we are labeling the comment for its stance toward the US Army as well as the video it is commenting on, this results in a total of 12 LLM labeling runs across the entire dataset.

### v3

@@ Abstract
Public opinion of military organizations significantly influences their ability to recruit talented individuals. As recruitment efforts increasingly extend into digital spaces like social media, it becomes essential to assess the stance of social media users toward online military content. However, there is a notable lack of data for analyzing opinions on military recruiting efforts online, compounded by challenges in stance labeling, which is crucial for understanding public perceptions. Despite the importance of stance analysis for successful online military recruitment, creating human-annotated, in-domain stance labels is resource-intensive. In this paper, we address both the challenges of stance labeling and the scarcity of data on public opinions of online military recruitment by introducing and releasing the DIVERSE dataset:  [^1] https://doi.org/10.5281/zenodo.10493803. This dataset comprises all comments from the U.S. Army’s official YouTube Channel videos. We employed a state-of-the-art weak supervision approach, leveraging large language models to label the stance of each comment toward its respective video and the U.S. Army. Our findings indicate that the U.S. Army’s videos began attracting a significant number of comments post-2021, with the stance distribution generally balanced among supportive, oppositional, and neutral comments, with a slight skew towards oppositional versus supportive comments.
@@ I Introduction
The U.S. military is facing a recruitment crisis. For the last few years, military organizations like the U.S. Army have struggled to meet recruiting targets to maintain their desired force structure (Brown, 2023). Social media plays an important role in discerning stances (AlDayel and Magdy, 2021), and in recruitment (Ng and Cruickshank, 2023). Thus, it is critical to understand how messages from the military are received on social media, as it can shape recruiting messaging. More broadly speaking, the ability to evaluate how social media content is being received by a target audience is crucial for any marketing campaign, be it a war recruiting effort or a organizational recruitment effort or for influencing consumer behavior.
@@
A particular indicator of the reception of social media content is stance. Stance detection, or the determination of an expressed or implied opinion toward a target, remains a primary task for many computational social studies. Several articles have proposed methods and datasets for the task of stance detection (Allaway and McKeown, 2023; Du Bois, 2007). However, existing datasets explore stance as expressed on a single social media platform, mainly Twitter/X, and only have a single stance per comment (Sobhani et al., 2017; Conforti et al., 2020; Mohammad et al., 2016; Hossain et al., 2020; Kochkina et al., 2018; Villa-Cox et al., 2020). Since stance detection remains a difficult task due to the context-dependent nature of the definition of stance (e.g., the requirement of a target for stance to be defined), along with varying definitions of stance (Ng and Carley, 2022), all of the aforementioned stance datasets rely on human annotation to create the stance labels. This can result in human errors being present in the datasets as well as greatly reduce the size of the datasets to be in the thousands since human annotation requires a lot of resources and time.
@@
To address these shortcomings in prior stance datasets, we propose a new stance dataset based on YouTube comments and labeled with machine-assisted labeling through the data programming paradigm of weak supervision. This dataset is a collection of the YouTube comments posted on videos uploaded by the U.S. Army [^2] Offical U.S. Army YouTube Channel: https://www.youtube.com/@usarmy. The comments are first annotated for weak labels representing aspects of stance expression (hate speech, sarcasm, sentiment). The comments are then labeled by prompting three different LLMs. Finally, these weak labels are combined to produce the final stance labels of the comments towards the U.S. Army and its videos. With this data set, we can better understand critical issues with the public opinion toward the U.S. Army and its recruitment efforts.
Our contributions in this paper are as follows:
@@
  • A novel benchmark stance dataset of YouTube comments. We present a dataset for the task of stance detection that has several novel differences from previous benchmark datasets. While most stance datasets use posts from Twitter, ours provides posts from the YouTube platform instead, contributing to the breadth of datasets. Another difference is that our dataset labels stances towards more than one target in a comment (the only other being Sobhani et al. (2017)), and provides labels to weak signals including the presence of contentious issues, conspiracy theories, and other emotional content. Finally, this dataset has more than three times the number of labeled comments over the current largest benchmark dataset (Conforti et al. (2020) at ${}32,000$ data points vs $>$216,000 data points in ours).
  • The use of a data-programming based stance-labeling methodology. We combine recent advances for creating large, high-quality labeled datasets with contextual knowledge to generate quality stance labels Ratner et al. (2019); Huang et al. (2024); Smith et al. (2024). This involves making use of weak labels such as the presence of hate speech, the sentiment of the content, the presence of sarcasm, known keywords, and the evaluation of Large Language Models (LLMs), which are combined using weak supervision to determine the final stance label.
@@ II Related Work
Finally, it is important to note that some parallel works to ours have explored combining weak supervision with LLMs to create labels. In Smith et al. (2024), the authors prompt an LLM to provide possible labels for examples and then combine these weak labels through weak supervision to produce final labels, which are significantly better than what an LLM can produce via zero-shot. Huang et al. (2024) use an LLM to create weak labeling functions for a dataset and then label the dataset with those weak labeling functions and, finally, combine the weak labels through weak supervision to produce a labeled dataset. In all cases, the combination of an LLM and weak supervision produces better improved labeling outcomes from a zero-shot scenario.
@@ III Dataset Creation
This dataset was collected to understand the reception of YouTube users toward the US Army’s marketing material on YouTube, thereby understanding their support/opposition towards the military. Each comment is labeled for two stances: (a) the stance of the comment towards its source video, and (b) the stance of the comment towards the US Army. There are three possible options for stances: “supports”, “against”, and “neutral”. The following parts of this section details the data collection and labeling procedures.
@@ III-A Data Collection
We collected comments on the videos of the US Army’s official YouTube channel (username: @usarmy [^3] https://www.youtube.com/usarmy) using the YouTube Data API [^4] https://developers.google.com/youtube/v3/docs/comments/list. This returned both comments as well as replies to comments. Collection took place from 2 to 5 October 2023 and again from 2-9 April 2024, where we retrieved a total of 216,385 comments across 1,082 videos. For each comment, we collected the following information:
@@
  • video_id: the unique id of the video of the comment
  • name: the title of the video of the comment
  • description: a text description of the video of the comment
@@ III-B Data Annotation Methodology
Instead of using human annotators for stance labeling, we used weak supervision Ratner et al. (2016, 2019). This paradigm produces a final, more accurate label using a series of lesser quality labels, weak labels, which may capture different aspects of the concept being labeled. Such labeling techniques have been employed in labeling happiness with cognitive appraisal dimension Liu and Jaidka (2023) and negotiation strategies using the presence of reasoning and friendliness Jaidka et al. (2023). We used a combination of weak labels based upon subject matter expertise, domain expertise, machine learning models, and recent advances in the use of LLMs for stance detection. These labels were then combined through a weak supervision model to produce the final labels. In this way, we include both specialty functions similar to Huang et al. (2024) as well as LLM-direct labels in Smith et al. (2024) into the weak supervision process. Figure 1, depicts the entire data annotation methodology workflow. We detail key aspects of the workflow in the following sections.
Figure 1: Pipeline of stance labeling methodology
@@ III-C Annotating Weak Labels
To create the weak labels we leveraged dataset-specific traits, domain expertise and LLMs. For example, in this dataset, the presence of hate speech or sarcasm frequently indicates the comment being against the video and by extension, the user that uploads it, which is the US Army. Therefore, the presence of these weak signals can be combined to indicate the final label. It is important to note for any given weak label, it need not produce a label for every comment in the dataset. Some of the weak labels simply do not apply to some of the comments, in which case they have an “abstain” response, which is distinct from the possible stance labels, and are not used in the computation of the final stance labels.
@@ III-C1 Hate Speech
Hate Speech is defined as speech to incite violence or hate, and typically has specific targets like ethnic groups or religion Fortuna and Nunes (2018). In the case of this dataset and stance, we noticed that many of the comments against both the posted videos and the US Army more generally featured aspects of hate speech, like antisemitic and misogynistic content. Several machine learning algorithms ranging from logistic regression to neural network models have been constructed to perform automatic hate speech detection Fortuna and Nunes (2018). For this weak label, we used the hate speech classification model from Kralj Novak et al. (2022)  [^5] https://huggingface.co/IMSyPP/hate_speech_en. This model was used as is without any fine-tuning. For each input text, this model returns three different types of “hate” label or an “acceptable” label. We considered the “hate” response as an “against” stance label, and any other response as an “abstain.”
@@ III-C2 Sarcasm
Sarcasm is a literary device to convey one’s intention in an implicit manner Chaudhari and Chandankhede (2017). As with hate speech, we found that commenters on these YouTube videos often use sarcasm as a way of attacking the video or the entity behind the video. We used a multilingual sarcasm detector666helinivan/multilingual-sarcasm-detector which trains a BERT-based sarcasm detector from newspapers text in English, Dutch and Italian. For each input text, this model returns a label of “sarcasm” or “no_sarcasm”, and we considered the “sarcasm” response as an “against” stance label, and any other response as an “abstain.”
@@ III-C3 Sentiment
Sentiment analysis provides an understanding of the positivity or negativity of public sentiment towards a topic Ibrohim et al. (2023). In the context of stance detection and classification, while stance is distinct from sentiment, the two often overlap; a negative sentiment often indicates a stance against a target and vice-versa for a positive sentiment. We used a multilingual-cased sentiment analysis model777lxyuan/distilbert-base-multilingual-cased-sentiments-student that performs sentiment-analysis using a teacher-student architecture with the DistillBERT model. For each input text, this model returns a score between 0 and 1 for three possible sentiment labels of “positive”, “negative”, and “neutral”. We considered a “positive” sentiment greater than 0.75 as a “supports” stance and a “negative” sentiment greater than 0.75 as a “against” stance and any other value as a “abstain” stance.
@@ III-C4 Use of Keywords
We manually determined a set of keywords and phrases that when present, would represent either one of the non-neutral stance labels. For each comment, we identified if any words from the set of keywords were present and assigned the corresponding stance to this weak label. For example, keywords and phrases like “hooah”, “awesome”, and “god bless” were associated with a “supports” stance, while “woke”, “murder”, and “propaganda” were associated with an “against” stance. the full list of these words is available in our released code [^8] https://doi.org/10.5281/zenodo.10493803.
@@ III-C5 Stance from Large Language Models
Due to the size of the dataset, we opted to use open-source LLMs which could be run on local hardware. Specifically, we used three different LLMs: (a) Flan-UL2 (Tay et al., 2022), which is an encoder-decoder architecture, (b) Mistral-7B (Jiang et al., 2023), which is a decoder-only architecture, and (c) Llama-3-8B Dubey et al. (2024), which is another decoder-only architecture. We used the Flan-UL2 and Llama-3 models in a zero-shot setting only (i.e., we did not perform any fine-tuning on the model) (Ziems et al., 2023). For the Mistral-7B model, we both used it in a zero-shot setting and we fine-tuned the model on a collection of benchmark stance datasets, including Semeval2016 (Mohammad et al., 2016), Election2016 (Sobhani et al., 2017), srq (Villa-Cox et al., 2020), wtwt (Conforti et al., 2020), phemerumors (Kochkina et al., 2018), and covid-lies (Hossain et al., 2020). Based on the research in (Ng and Carley, 2022), which found that the inclusion of more benchmark datasets into training improved the generalizability of the trained model, hence we trained on all of the datasets. For the training of the Mistral LLM, we used a LoRA adapter (Hu et al., 2021) with parameters of $\alpha=1$, $r=16$, $dropout=0.1$, and no bias term. The model was trained for two epochs using an AdamW optimizer (Loshchilov and Hutter, 2017), a learning rate of $41e-$ 4, and a linear learning rate scheduler with 50 warm-up steps. To feed the data to the model during training, we adopted a prompt that included context about stance classification, as has been done in zero-shot prompting for stance in other works. The following was the training prompt:
@@
Where the [dataset entity] was varied depending on the benchmark dataset (e.g., for Election2016, it was ‘politician’, while for phemerumors it was ‘rumor’). For the srq dataset, we changed the prompt slightly to incorporate the post that the reply is replying to as the task in that dataset is to label the stance of the reply to the post, versus the stance of the post about a topic or person. The loss was determined by whether the model could output the tokens for the correct stance, and only those tokens. The training of the model was done on a local computer running Ubuntu Linux with a 96-core processor and three A6000 GPUs. Overall, this resulted in four, different LLMs for labeling the comments: one off-the-shelf UL2 model, one off-the-shelf Llama-3 model and two Mistral models.
Finally, we also fine-tuned a small BERT model as was done in Ng and Carley (2022) on all of the benchmark datasets previously described and labeled the stances with that fine-tuned model as well.
@@
In analyzing the comments for stances using LLMs, we used two different prompts, one that only included the comment, and one that included the comment that a comment was replying to, if it was replying to a comment. The labeling prompt without incorporating the reply structure is:
@@
We also experimented with a question answering phrasing of the task during labeling, but found the analyze phrasing to generally perform slightly better in benchmark datasets. Overall, each comment was labeled for a stance by an LLM seven different times: once with the reply prompt and once without the reply prompt for the UL2 model, the Llama-3 model, and other Mistral model, and with the without the reply prompt by the stance-tuned Mistral model. Since we are labeling the comment for its stance toward the US Army as well as the video it is commenting on, this results in a total of 14 LLM labeling runs across the entire dataset.
To combine the LLM-derived labels we choose to ensemble the LLM labels together using majority vote. Specifically, we considered those labels where more than half (i.e., 4 or more of the LLM labelers) agreed on a stance for a comment as producing a stance, otherwise we assigned the label from the LLMs as “abstain.” Although this minorly reduced some of the coverage of the LLM labels, it also improved the quality of the LLM labels overall.